OpenStudy (anonymous):

Simplify completely: -3(4-9)+6-2(-8+5)

pooja195 (pooja195):

ok frist you have to do wat?

pooja195 (pooja195):

parenthses rite so wat is 4-9

pooja195 (pooja195):

u there

OpenStudy (anonymous):

Im here

pooja195 (pooja195):

ok good 4-9

OpenStudy (anonymous):

The Answer is 27

pooja195 (pooja195):

ok yr rite

OpenStudy (anonymous):

1.) Distribute -3(4-9) =-12+27 2.) In between the two distributing factors 6 Note: Leave this number alone until the end Distribute -2(8+5) = -16-10 Now add/subtract them all together -12+27+6-16-10= ?

OpenStudy (anonymous):

Im here